OpenStudy (anonymous):

Find the indicated missing measurement. Include the second solution if applicable. What am I doing wrong?

OpenStudy (anonymous):

 So this is what I have so far: sinA/a = sinB/b sin34 degrees/9 = sinB/7 sinB = 7sin34 degrees/9 sinB = 25.8 degrees... Angle A is less than 90 degrees -> a is greater than or equal to bsinA bsinA = 7sin34 = 3.91 a is greater than or equal to bsinA = 2 solutions B' = 180-25.8 = 154.2 degrees But when I find for angle C of the second triangle it doesn't add up to 180... What am I doing wrong?

OpenStudy (mertsj):

\[\frac{\sin 34}{9}=\frac{\sin B}{7}\] \[\sin B=\frac{7\sin 34}{9}=.4349278\] B=25.78

OpenStudy (mertsj):

C=180-25.78-34=120.22

OpenStudy (anonymous):

don't I have to subtract 25.8 from 180 to get the angle of B'?

OpenStudy (mertsj):

There is nothing ambiguous about this problem because of the given information. You know that angle B is less than angle A because the side opposite angle B is less than the side opposite angle A. After you find A, use the fact that the sum of the angles of a triangle is 180

OpenStudy (kropot72):

sin B = (7 * sin 34 degrees)/9 = 0.4349 \[B=\sin^{-1} 0.4349=25.8degrees\] C = 180 degrees - (25.8 + 34) = 120.2 degrees \[\frac{a}{\sin A}=\frac{c}{\sin C}\] \[\frac{9}{\sin A}=\frac{c}{\sin 120.2}\] c = 13.9

OpenStudy (mertsj):

The second solution is not applicable.

OpenStudy (mertsj):

In other words, there is no B' in this problem.

OpenStudy (anonymous):

ohhh.. thank you guys!!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

but.. wait.

OpenStudy (anonymous):

when do I use sine law ambiguous case then?

OpenStudy (anonymous):

because isn't angle A an acute angle? Therefore a>b = 1 ohh..

OpenStudy (anonymous):

a is greater than b... which means one solution.